A processor makes two components, A and B, which are then packaged together as the final product (each product sold contains one A and one B). The processor can do only one component at a time: either it can make As or it can make Bs. There is a setup time when switching from A to B.

Current plans are to make 100 units of component A, then 100 units of component B, then 100 units of component A, then 100 units of component B, and so forth, where the setup and run times for each component are given below.

COMPONENT SETUP/CHANGEOVER TIME RUN TIME/UNIT
A 5 minutes 0.3 minute
B 10 minutes 0.1 minute

Assume the packaging of the two components is totally automated and takes only two seconds per unit of the final product. This packaging time is small enough that you can ignore it.

What is the average hourly output, in terms of the number of units of packaged product (which includes one component A and one component B)? (Round
Answer #1

Answer: - According to given data

Inorder to calculate the average hourly rate we need to find the opperation time of component of A and B

For component A = setup time + run time

Where setup time = 5 minutes

run time of 100 units of component A = 100 * 0.3 minutes = 30

Operation time of A = 5 + 30 = 35 minutes

For component B

Where setup time = 10 minutes

run time of 100 units of component B = 100 * 0.10 = 10

Operation time of B = 10 + 10 = 20 minutes

Average hourly rate is obtained the combination of two operations time to produce both the components of A and B of each

= 35 minutes + 20 minutes = 55 minutes per 100 units

For 60 minutes = 60*100 / 55 = 109 units

So Average hourly output in terms of units = 109 units